# Hierarchical Generative Adversarial Imitation Learning with Mid-level Input Generation for Autonomous Driving on Urban Environments

### v4

@@ VI Conclusion
Both GAIL and CGAN in hGAIL learns simultaneously to control the agent and generate the input representations, respectively. The learning takes place in an urban city without pedestrians or other cars, but with dynamic routes that can change the path on the fly. Our experiments have shown that the mid-level input generated by CGAN is essential for the learning task as the GAIL exclusively from cameras (without BEV) fails to even learn the task, keeping a high-infraction rate through training. The BC agent with its associated GAN can complete some turns in the trajectories, but not consistently as hGAIL can. In fact, hGAIL, after training exclusively on ten expert trajectories from one city, was able to generalize its navigation skills to a new, unseen city, and complete 98% of all 48 possible turns in the eight T intersections from that new city.
This work has demonstrated the usefulness of mid-level BEV input for realistic navigation scenarios, but also that this input representation can be learned concomitantly with the agent’s policy training. Thus, the BEV generation is learned with the same data distribution used to train the agent’s policy, i.e., the expert dataset and the data generated by the agent . In future work, we are interested in: adding dynamic obstacles such as pedestrians and other vehicles, as well as traffic lights and other weather conditions; testing the generalization of the policy to other new cities; and sim2real experiments .

### v5

@@ VI Conclusion
Both GAIL and CGAN in hGAIL learns simultaneously to control the agent and generate the input representations, respectively. The learning takes place in an urban city initially without pedestrians or other cars, but with dynamic routes that can change the path on the fly. In the following experiment, hGAIL is extended so that the agent can learn to respect traffic lights and deal with the detection of pedestrians and vehicles. Our experiments have shown that the mid-level input generated by CGAN is essential for the learning task as the GAIL exclusively from cameras (without BEV) fails to even learn the task, keeping a high-infraction rate through training. The BC agent with its associated GAN can complete some turns in the trajectories, but not consistently as hGAIL can. In fact, hGAIL, after training exclusively on ten expert trajectories from one city, was able to generalize its navigation skills to a new, unseen city, and complete 98% of all 48 possible turns in the eight T intersections from that new city.
This work has demonstrated the usefulness of mid-level BEV input for realistic navigation scenarios, but also that this input representation can be learned concomitantly with the agent’s policy training. Thus, the BEV generation is learned with the same data distribution used to train the agent’s policy, i.e., the expert dataset and the data generated by the agent. In future work, we are interested in extending the method making the policy generalize to other weather conditions and other new cities; and sim2real experiments. Also, we would like to investigate more closely the rarer situations that lead to infractions, proposing a more refined architecture to handle these difficult cases, as well as improve the quality of the expert.
APPENDIX
VI-A Generative Adversarial Imitation Learning (GAIL)
VI-B CGAN Network Architecture
Both CGAN networks’ architectures are presented in Fig. 10, where the common layers in orange refer to layers existing in both the Generator and the Discriminator. Notice that they are separate networks which do not share parameters: the figure was made to not repeat equivalent layers when describing both networks.
Generator
Discriminator
VI-C GAIL policy architecture
VI-D Hyperparameters
Tables  III and Table IV show the hyperparameter values for the GAIL and GAN parts of hGAIL, respectively .
TABLE III: Hyperparameters for GAIL
Description Value
Parallel environments ($N$) $6$
Initial adam step size (lr) $2.0\times 10^{-5}$
Adam step size exponential decay ($\lambda_{lr}$) $0.96$
Number of PPO epochs (K) $20$
Mini-batch size (m) $256$
Discount ($\gamma$) $0.99$
GAE parameter ($\lambda$) $0.9$
Clipping parameter ($\epsilon$) $0.2$
Value Function clipping parameter ($\epsilon_{vf}$) $0.2$
Value Function coefficient ($c_{1}$) $0.5$
Entropy coefficient ($\lambda_{\mathrm{ent}}$) $0.01$
Exploration coefficient ($\lambda_{\mathrm{exp}}$) $0.05$
Timesteps per epoch (T) $12288$
Weight of BC in GAIL loss ($\alpha$) $0.004$
Discriminator adam step size (lr) $2.5\times 10^{-4}$
Number discriminator epochs ($J$) $2$
TABLE IV: Hyperparameters for CGAN
Description Value
Adam step size (lr) $2.0\times 10^{-4}$
Number of GAN epochs (L) $4$
Mini-batch size (m) $32$
Patch size ($\gamma$) $(10,10)$
Resize ($\lambda$) $(192,192)$
Lambda pixel ($\epsilon$) $100$
VI-E Training progress
VI-F Evaluation environment
Town02 in Fig. 13 is the environment used for testing the generalization capabilities of the trained agents.
Figure 13: Town02 evaluation city , with one of the routes used to test agents trained in town01. The highlighted path has 1010 meters, 29 points in the sparse trajectory (shown as yellow dots) and 1030 points in the dense point trajectory (not shown).
ACKNOWLEDGMENT
The authors acknowledge CAPES (Coordination for the Improvement of Higher Education Personnel) - Brazil for financial support.